Formulate at least two problems invovlimg arcs and central angles,thensolve


1
Expert's answer
2021-01-26T01:15:46-0500

1.Find m∠CAE


The degree measure of a minor arc is equal to the measure of the central angle that intercepts it. Therefore,

m∠CAE=140°


2.Calculate the arc length of a sector whose central angle is 45° and the radius of 12 cm.


Arc Length ="\frac{\theta} {360}2\pi r"


="\frac{45}{360}×2×3.142×12"


=9.426cm